OpenStudy (thebadyboy):

For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (thebadyboy):

@King.Void.

OpenStudy (ilovecake):

A

OpenStudy (thebadyboy):

there can be more than 1

OpenStudy (ilovecake):

I think that's the only one

OpenStudy (thebadyboy):

ok thx i have more

OpenStudy (thebadyboy):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (freckles):

are all of these products in the set {0,1} 0*1 1*0 1*1 0*0 ---- are all these quotients in the set {0,1} 0/1 1/0 1/1 0/0 (hint here: can you divide by 0?) ---- are all of these sums in the set {0,1} 0+1 1+0 1+1 0+0 ----- are all of these differences in the set {0,1} 0-1 1-0 1-1 0-0 ---- you are looking for a yes the question you say yes to is the operation in which the set {0,1} is closed under

OpenStudy (thebadyboy):

so a

OpenStudy (freckles):

whole numbers are {0,1,2,3,4,5,...} so which one shows adding two whole numbers and it producing a whole number

OpenStudy (thebadyboy):

ok

OpenStudy (ilovecake):

He is saying it's A. As for the second one it would be C and D. Good Luck.

OpenStudy (ilovecake):

Would you like explanation for the first one?